Direction: Using the information given on each item, write the

standard form equation of each ellipse.

1. Vertices: (12, -4), (-14, -4)

Foci: (11, -4), (-13, -4)

2. Vertices: (6,22), (6, -4)

Foci: (6, 14), (6,4)

3. Foci: (0,9), (-10,9)

Co-vertices: (-5, -21), (-5, -3)

4. Foci: (5,-8), (5,-8)

Co-vertices: (0,4), (0, -20) 5. Center: (6, 2)

Vertex: (6, -4) Co-vertex: (9,2)


1
Expert's answer
2021-11-12T15:40:48-0500

1.

The y-coordinates of the vertices and foci are the same, so the major axis is parallel to the x-axis. Thus, the equation of the ellipse will have the form;

"\dfrac{{\left(x-k\right)}^{2}}{{a}^{2}}+\dfrac{{\left(y-h\right)}^{2}}{{b}^{2}}=1"


First, we identify the center. The center is halfway between the vertices, (12, -4), (-14, -4). Applying the midpoint formula, we have:

"\begin{aligned}\left(k,h\right)&=\left(\dfrac{12+\left(-14\right)}{2},\dfrac{-4+\left(-4\right)}{2}\right) \\ &=\left(-1,-4\right) \end{aligned}"


Next, we find a². The length of the major axis, 2a, is bounded by the vertices. We solve for a by finding the distance between the x-coordinates of the vertices.

"\begin{aligned}2a&=-14-\left(12\right)\\ 2a&=-26\\ a&=-13\end{aligned}"

a² = 169


Now we find c². The foci are given by (k±c, h). So, (k−c, h) = (11, −4) and (k+c, h) = (−13, -4). We substitute k = −3 using either of these points to solve for c.

"k+c = -13\\
-1+ c = -13\\
c = -12"

c² = 144


Next, we solve for b² using the equation c² = a² − b².

"c² = a²-b²\\
144 = 169- b²\\
b² = 25\\
b= 5"


"\dfrac{{\left(x+1\right)}^{2}}{{169}}+\dfrac{{\left(y+4\right)}^{2}}{{25}}=1"




2.

The x-coordinates of the vertices and foci are the same, so the major axis is parallel to the y-axis. Thus, the equation of the ellipse will have the form;

"\dfrac{{\left(x-h\right)}^{2}}{{b}^{2}}+\dfrac{{\left(y-k\right)}^{2}}{{a}^{2}}=1"


(h, k) which is the midpoint is;

(h, k) = (6, 9)


Next, we find a². The length of the major axis, 2a, is bounded by the vertices. We solve for a by finding the distance between the y-coordinates of the vertices.

"2a = -4-22 \\
2a = -26\\
a = -13\\
a² = 169"


"k + c = 4\\
9+c = 4\\
c = -5\\
c² = 25"


"c² = a²- b²\\
25 = 169 -b²\\
b² = 144\\
b = 12"


"\dfrac{{\left(x-6\right)}^{2}}{{144}}+\dfrac{{\left(y-9\right)}^{2}}{{169}}=1"



3.

The equation of an ellipse is "\dfrac{(x-h)^2}{a^2} +\dfrac{(y-k)^2}{b^2}=1" for a horizontally oriented ellipse and "\dfrac{(x-h)^2}{b^2} +\dfrac{(y-k)^2}{a^2} =1" for a vertically oriented ellipse.

where (h, k) is the center and the distance c from the center to the foci is given by "a^2-b^2=c^2". a is the distance from the center to the vertices and b is the distance from the center to the co-vertices.


The center of the ellipse is half way between the vertices. Thus, the center (h, k) of the ellipse is (0, 0) and the ellipse is vertically oriented.

a is the distance between the center and the vertices, so a = c is the distance between the center and the foci, so c = 4


"\dfrac{(x-9)^2}{20^2} +\dfrac{(y-12)^2}{5^2} =1"


4.

The equation of an ellipse is"\dfrac{(x-h)^2}{b^2} +\dfrac{(y-k)^2}{a^2} =1"

where (h, k) is the center and the distance c from the center to the foci is given by "a\u00b2+b\u00b2= c\u00b2". a is the distance from the center to the vertices and b is the distance from the center to the co-vertices.

The center of the ellipse is half way between the vertices. Thus, the center (h, k) of the ellipse is (0, 0) and the ellipse is vertically oriented.

i.e h = 0, k = 0

a which is the x axis point is 4, while b is 20

The equation becomes

"\dfrac{x^2}{4^2} +\dfrac{y^2}{20^2} =1"



5.




Center: (6, 2) Vertex: (6, -4) Co-vertex: (9,2)


from the image above,

vertices = (6, -4), (6, 8) => -4+12

covertices = (9, 2) | (3, 2) => 9-3-3

semimajor axis length = 6

semiminor axis length = 6/3 = 2

h = 6

k = 2

"a² -b² = c²\\
6² -b² = 5²\\
b² = 9\\ 
b = 3"


"\therefore" the equation becomes;


"\dfrac{(x - 6)^2 }9+ \dfrac{(y - 2)^2) }{36}= 1"